# Identifying Query-Relevant Neurons in Large Language Models for Long-Form Texts

### v3

@@ Abstract
Large Language Models (LLMs) possess vast amounts of knowledge within their parameters, prompting research into methods for locating and editing this knowledge. Previous work has largely focused on locating entity-related (often single-token) facts in smaller models. However, several key questions remain unanswered: (1) How can we effectively locate query-relevant neurons in contemporary autoregressive LLMs, such as Llama and Mistral? (2) How can we address the challenge of long-form text generation? (3) Are there localized knowledge regions in LLMs? In this study, we introduce Query-Relevant Neuron Cluster Attribution (QRNCA), a novel architecture-agnostic framework capable of identifying query-relevant neurons in LLMs. QRNCA allows for the examination of long-form answers beyond triplet facts by employing the proxy task of multi-choice question answering. To evaluate the effectiveness of our detected neurons, we build two multi-choice QA datasets spanning diverse domains and languages. Empirical evaluations demonstrate that our method outperforms baseline methods significantly. Further, analysis of neuron distributions reveals the presence of visible localized regions, particularly within different domains. Finally, we show potential applications of our detected neurons in knowledge editing and neuron-based prediction.
@@ 1 Introduction
Large Language Models (LLMs) contain substantial amounts of knowledge within their neurons (or parameters). Recent research has focused on identifying and localizing these knowledge neurons to gain insights into the information processing mechanisms of LLMs. Activation-based methods (Voita, Ferrando, and Nalmpantis 2023) examine activation patterns to elucidate the role of neurons in the reasoning process. However, these methods often struggle to directly attribute specific outputs to corresponding inputs, thereby limiting their effectiveness in accurately identifying relevant knowledge. Gradient-based methods (Dai et al. 2022) measure the sensitivity of model outputs to internal components in response to specific inputs, which enables the effective identification of neurons relevant to particular queries. However, these methods typically employ fill-in-the-blank tasks, such as ‘‘Paris is the capital of  ’’, to ascertain correlations of triplet facts. Causality-based methods (Meng et al. 2022a) take a different approach by employing causal mediation analysis to pinpoint layers within LLMs that store factual associations. Another branch of pioneering research attempts to locate functional regions in small-size language models such as BERT (Kenton and Toutanova 2019) and GPT-small (Radford et al. 2019), including linguistic regions (Zhang et al. 2024b), factual subnetworks (Ren and Zhu 2022; Bayazit et al. 2023), and modular structures (Zhang et al. 2023; Conmy et al. 2023).
While these studies successfully identify knowledge associations stored within LLMs, three significant questions remain underexplored: (1) How can we effectively locate query-relevant neurons in contemporary autoregressive LLMs, such as Llama (Touvron et al. 2023) and Mistral (Jiang et al. 2023), given their large model size and different architectures? (2) How can we address the challenge of long-form text generation, as previous methods have been limited to triplet facts? (3) Are there localized knowledge regions in LLMs analogous to the localized functional regions observed in human brains (Brett, Johnsrude, and Owen 2002)?
@@
To address the first two questions, we introduce a novel framework named Query-Relevant Neuron Cluster Attribution (QRNCA) designed to identify query-relevant neurons in LLMs. The principal advantages of our framework are its architecture-agnostic nature and its capability of handling long-form text generation effectively, as shown in Table 1. QRNCA aims to extract Query-Relevant (QR) neurons for each input query. The process begins by transforming an open-ended generation task into a multiple-choice question-answering format. By employing prompt engineering, we constrain LLMs to generate only the option letter rather than the complete answer. This approach allows for the examination of long-form generation beyond single tokens. Subsequently, we adapt the Knowledge Attribution method (Dai et al. 2022) to compute Neuron Attribution, which elucidates the relationship between neurons and the input query. We then gather clusters for a series of queries and calculate the Inverse Cluster Attribution. This step mitigates the influence of neurons that recur across clusters (or queries). The final step involves multiplying the neuron attribution and inverse cluster attribution values to pinpoint correlated neurons. Additionally, we identify certain Common Neurons that are associated with common words, punctuation marks, and option letters. Excluding these common neurons enhances the detection of QR neurons. Empirical evaluations demonstrate that our proposed method outperforms baseline approaches.
@@ 2.1 Locating Knowledge in LLMs
Large Language Models contain a vast range of knowledge within their parameters, spanning factual (Petroni et al. 2019; Zhou et al. 2020; Jiang et al. 2020; Roberts, Raffel, and Shazeer 2020; Pezeshkpour 2023), linguistic (Liu et al. 2019; Jawahar, Sagot, and Seddah 2019; Chen, Varoquaux, and Suchanek 2023), and domain-specific information (Sung et al. 2021; Frieder et al. 2024). Recent mechanistic studies suggest that knowledge is primarily stored in the FFNs (Feed-forward Networks) layers of Transformers (Geva et al. 2021, 2022), which prompts ongoing research efforts aimed at developing methods to identify and locate this knowledge within these layers. Activation-based methods (Voita, Ferrando, and Nalmpantis 2023; Gurnee et al. 2024) investigate the activation patterns of neurons to interpret how the network processes information at different stages. However, a key limitation of these methods is their inability to directly attribute the model’s output to specific inputs, which limits their precision in identifying relevant knowledge. Gradient-based methods (Ancona et al. 2019; Dai et al. 2022), on the other hand, offer fine-grained attribution by quantifying the sensitivity of model outputs to internal components in response to a given input. This approach effectively identifies neurons relevant to specific queries. Nonetheless, current gradient-based techniques have primarily focused on single-token triplet facts. Another approach, Causality-based methods, employs causal mediation analysis to discern the particular layers associated with a given factual input (Meng et al. 2022a). This line of research has evolved into a locate-and-edit paradigm, aimed at refining knowledge within LLMs (Meng et al. 2022b; Ju and Zhang 2023; Zhang et al. 2024a). In addition to general knowledge locating approaches, recent studies have focused on identifying neurons responsible for specific tasks, such as linguistic (Chen et al. 2024b; Tang et al. 2024; Kojima et al. 2024), privacy-related (Wu et al. 2023; Chen et al. 2024a) and bias-related neurons (Yang, Kang, and Jung 2023).
In this work, we propose a novel gradient-based attribution method aimed at locating input-output knowledge within LLMs. Unlike previous methodologies, our approach mainly focuses on long-form texts beyond entity facts.
@@ 3 Background
Dai et al. (2022) propose a gradient-based Knowledge Attribution to identify the knowledge neurons in BERT by using the fill-in-the-blank cloze task. Their method evaluates the contribution of each neuron in FFNs to the knowledge predictions. Given a prompt $q$ (‘‘Paris is the capital of  ’’), the probability of the correct answer predicted by an LLM can be formulated as:
@@
where $m$ is the number of the estimation steps. Finally, they identify a coarse set of knowledge neurons whose attribution scores are greater than a threshold $t$.
4 Locating Query-Relevant (QR) Neurons in Autoregressive LLMs
@@
To clarify the main concepts in our framework, we provide the following key notions: QR Neuron is an individual neuron highly correlated with a specific query, capable of influencing the corresponding knowledge expression. Our primary objective is to identify a set of QR neurons for a given input. QR Cluster represents a coarse grouping of neurons associated with a specific query. This cluster may include noisy neurons and require further refinement. Common Neuron is consistently activated by a wide range of inputs, representing general knowledge or concepts.
The overall framework is shown in Figure 1. Our framework first resorts to the proxy task of Multi-Choice QA to deal with long-form texts. Starting with a given query, the framework employs Neuron Attribution to derive a QR Cluster. Each neuron within this cluster is assigned an attribution score that indicates its relevance to the query. Next, we apply Inverse Cluster Attribution to attenuate the influence of neurons that frequently occur across multiple clusters. Finally, we identify Common Neurons, as those lacking discriminative power in determining query relevance and representing common knowledge or concepts. Refining the extraction of QR neurons by excluding these common neurons enhances the precision in identifying critical neural correlates.
@@ 4.1 Multi-Choice QA Transformation
Multi-choice QA is widely used in a variety of real-world educational assessments and standardized tests. Meanwhile, many known benchmarks such as MMLU (Hendrycks et al. 2020) and Big-bench (Srivastava et al. 2023) use multi-choice QA to evaluate the breadth and depth of a model’s knowledge. Therefore, we adopt the proxy task of multi-choice QA to study the knowledge associations in LLMs. Given the biology question “The energy given up by electrons as they move through the electron transport chain is used to?”, the correct answer can be the long-form text “produce ATP”. To deal with long-form answers, we advocate for the transformation of questions and their corresponding answers into a multiple-choice framework, as illustrated in Figure 1. This approach involves the generation of incorrect options by randomly sampling answers within the same domain. Following this, the LLM is prompted to produce only the option letter. Subsequently, we investigate the neurons correlated with the input query. To mitigate the impact of randomness, we devise multiple prompt templates and systematically shuffle the order of options to prevent the model from learning spurious correlations based on option letters. These prompt templates are detailed in Table A3 in the Supplementary Material (SM).
@@ 4.2 Neuron Attribution
Here, the numerator means that we sum up the scores of different instantiated templates together as the initial attribution score. The denominator $Z$ is the normalized factor by summing the initial attribution scores of all neurons. Since the number of prompts for each query may be varied and the initial attribution scores may be scaled differently, we use normalization to make the attribution scores comparable across queries.
@@ 4.4 Common Neurons
We observe that some neurons with a relatively high attribution score are still shared across clusters. Through case studies (as shown in Table 4), we demonstrate that they express commonly used concepts such as option letters (‘‘A’’ and ‘‘B’’) or stop words (‘‘and’’ and ‘‘the’’). Therefore, we count the frequency of each neuron across clusters. If the frequency is higher than the $u\%$ of total clusters, we assign the given neuron into the common neuron set.
@@ 4.5 Obtaining QR Neurons
Given a query, the final score of a neuron can be computed as:
@@ 5.1 Experimental Settings
To mitigate sensitivity to prompts and option orders, each query is instantiated with multiple distinct templates (as shown in Table A3 in the SM), and the option orders are shuffled each time. The statistics of our datasets are shown in Table 2 and examples can be found in Table A4 in the SM.
@@
We compare our QRNCA to other neuron-level baselines [^1] We do not compare to ROME (Meng et al. 2022a) since it locates layers instead of neurons. Also, we do not compare to task-specific methods.: Random Neuron are randomly selected from FFNs and we make sure they have the same number of neurons as QRNCA; Activation selects neurons with high activated values. Kowledge Neuron is adapted from knowledge attribution (Dai et al. 2022) by using the multi-choice QA task; QRNCA wo/ ICA only uses neuron attribution (Eq 5) to obtain relevant neurons, which dose not involve the computation of Inverse Cluster Attribution; QRNCA w/ Common Neuron is a variant without removing common neurons.
@@
Table 3: Comparisons of different knowledge locating methods for Llama-2-7B. The metric here is the Probability Change Ratio (PCR) described in Section 5.1. Details are shown in Table A2 in the appendix.
@@
As for the used hyper-parameters, the number of estimation steps was set to $m=16$ and the attribution threshold $t$ to 0.2 times the maximum attribution score. The template number was $|\mathcal{Q}|=3$, the frequency $u$ for obtaining common neurons was 30%, and the top-$v$ for select coarse neurons was 20. We ran all experiments on three NVIDIA-V100. It took 120 seconds on average to locate neurons for a query with three prompt templates. For each domain and language, the average number of detected QR neurons is between 12 and 17 (as shown in Table A1 in the SM).
@@ 5.3 QR Neurons Can Impact the Knowledge Expression
Table 3 presents the overall performance of various methods. Our QRNCA method consistently outperforms other baselines, evidenced by its higher PCR. This indicates that our identified QR neurons significantly affect the probability of correct answers while exerting a relatively low impact on unrelated queries. For instance, our method achieves a boosting ratio of 41.2 on the language dataset, the highest among the baselines. Additionally, both our proposed ICA and the removal of common neurons can bring benefits to locating neurons, which is validated by the worse performance of the two QRNCA variants.
Furthermore, Figure 3 illustrates the percentage change in probability for each domain and language after boosting neuron values. Again, we can clearly observe the effectiveness of our detected QR neurons. Additionally, we performed supplementary experiments on Mistral-7B. The results, presented in Figure A4 in the SM, consistently support our conclusions.
@@ 7 Conclusion
In this study, we introduce a novel framework, QRNCA, for identifying query-relevant neurons in contemporary autoregressive language models. QRNCA leverages a multi-choice QA proxy task to address the complexity of long-form answers, extending beyond triplet facts. Meanwhile, it adopts strategies of inverse cluster attribution and common neuron removal to refine QR neurons. To validate our approach, we curate two datasets encompassing diverse domains and languages. Our experimental results show that our method outperforms existing baselines in identifying associated neurons. Additionally, this study pioneers the exploration of localized knowledge regions in LLMs and demonstrates Llama contains knowledge-specific regions in the middle layers while language-specific neurons tend to be distributed across different layers. Further, we prototype two potential usages of identified neurons in applications such as knowledge editing and neuron-based prediction. We hope that our findings are beneficial for further research in understanding the knowledge mechanisms underlying LLMs.

### v4

@@ Abstract
[…]
@@ 1 Introduction
Large Language Models (LLMs) contain substantial amounts of knowledge within their neurons (or parameters). Recent research has focused on identifying and localizing these knowledge neurons to gain insights into the information processing mechanisms of LLMs. Activation-based methods (Voita, Ferrando, and Nalmpantis 2023) examine activation patterns to elucidate the role of neurons in the reasoning process. However, these methods often struggle to directly attribute specific outputs to corresponding inputs, thereby limiting their effectiveness in accurately identifying relevant knowledge. Gradient-based methods (Dai et al. 2022) measure the sensitivity of model outputs to internal components in response to specific inputs, which enables the effective identification of neurons relevant to particular queries. However, these methods typically employ fill-in-the-blank tasks, such as “Paris is the capital of  ”, to localise components representing triplet facts. Causality-based methods (Meng et al. 2022a) take a different approach by employing causal mediation analysis to pinpoint layers within LLMs that store factual associations. Another branch of pioneering research attempts to locate functional regions in small-size language models such as BERT (Kenton and Toutanova 2019) and GPT-small (Radford et al. 2019), including linguistic regions (Zhang et al. 2024b), factual subnetworks (Ren and Zhu 2022; Bayazit et al. 2023), and modular structures (Zhang et al. 2023; Conmy et al. 2023).
While these studies successfully identify knowledge associations stored within LLMs, three significant questions remain underexplored: (1) How can we effectively locate query-relevant neurons in contemporary decoder-only LLMs, such as Llama (Touvron et al. 2023) and Mistral (Jiang et al. 2023), given their large model size and different architectures? (2) How can we address the challenge of long-form text generation, as previous methods have been limited to triplet facts? (3) Are there localized knowledge regions in LLMs analogous to the localized functional regions observed in human brains (Brett, Johnsrude, and Owen 2002)?
@@
To address the first two questions, we introduce a novel framework named Query-Relevant Neuron Cluster Attribution (QRNCA) designed to identify query-relevant neurons in LLMs. The principal advantages of our framework are its architecture-agnostic nature and its capability of handling long-form text generation effectively, as shown in Table 1. QRNCA aims to extract Query-Relevant (QR) neurons for each query-answer fact. The process begins by transforming a free-form generation task into a multiple-choice question-answering format. By employing prompt engineering, we constrain LLMs to generate only the option letter rather than the complete answer. This approach allows for the examination of long-form generation beyond single tokens. Subsequently, we adapt the Knowledge Attribution method (Dai et al. 2022) to compute Neuron Attribution, which elucidates the relationship between neurons and the factual knowledge. We then gather clusters for a series of queries and calculate the Inverse Cluster Attribution. This step mitigates the influence of neurons that recur across clusters (or queries). The final step involves multiplying the neuron attribution and inverse cluster attribution values to pinpoint correlated neurons. Additionally, we identify certain Common Neurons that are associated with common words, punctuation marks, and option letters. Excluding these common neurons enhances the detection of QR neurons. Empirical evaluations demonstrate that our proposed method outperforms baseline approaches.
@@ 2.1 Locating Knowledge in LLMs
Large Language Models contain a vast range of knowledge within their parameters, spanning factual (Petroni et al. 2019; Zhou et al. 2020; Jiang et al. 2020; Roberts, Raffel, and Shazeer 2020; Pezeshkpour 2023), linguistic (Liu et al. 2019; Jawahar, Sagot, and Seddah 2019; Chen, Varoquaux, and Suchanek 2023), and domain-specific information (Sung et al. 2021; Frieder et al. 2024). Recent mechanistic studies suggest that knowledge is primarily stored in the FFN (Feed-forward Network) layers of Transformers (Geva et al. 2021, 2022), which prompts ongoing research efforts aimed at developing methods to identify and locate this knowledge within these layers. Activation-based methods (Voita, Ferrando, and Nalmpantis 2023; Gurnee et al. 2024) investigate the activation patterns of neurons to interpret how the network processes information at different stages. However, a key limitation of these methods is their inability to directly attribute the model’s output to specific inputs, which limits their precision in identifying relevant knowledge. Gradient-based methods (Ancona et al. 2019; Dai et al. 2022), on the other hand, offer fine-grained attribution by quantifying the sensitivity of model outputs to internal components in response to a given input. This approach effectively identifies neurons relevant to specific queries. Nonetheless, current gradient-based techniques have primarily focused on single-token triplet facts. Another approach, Causality-based methods, employs causal mediation analysis to discern the particular layers associated with a given factual input (Meng et al. 2022a). This line of research has evolved into a locate-and-edit paradigm, aimed at refining knowledge within LLMs (Meng et al. 2022b; Ju and Zhang 2023; Zhang et al. 2024a). In addition to general knowledge locating approaches, recent studies have focused on identifying neurons responsible for specific tasks, such as linguistic (Chen et al. 2024b; Tang et al. 2024; Kojima et al. 2024), privacy-related (Wu et al. 2023; Chen et al. 2024a) and bias-related neurons (Yang, Kang, and Jung 2023).
In this work, we propose a novel gradient-based attribution method aimed at locating input-output knowledge within LLMs. Unlike previous methodologies, our approach mainly focuses on long-form (or free-form) texts beyond entity facts.
@@ 3 Background
Dai et al. (2022) propose a gradient-based Knowledge Attribution to identify the knowledge neurons in BERT by using the fill-in-the-blank cloze task. Their method evaluates the contribution of each neuron in FFNs to the knowledge predictions. Given a query prompt $q$ (“Paris is the capital of  ”), the probability of the correct answer predicted by an LLM can be formulated as:
@@
where $m$ is the number of the estimation steps. Finally, they identify a coarse set of knowledge neurons whose attribution scores are greater than a threshold $t$. The localized neurons are supposed to be highly associated with a piece of knowledge, i.e., the query-answer facts.
4 Locating Query-Relevant (QR) Neurons in Decoder-only LLMs
@@
To clarify the main concepts in our framework, we provide the following key notions: QR Neuron is an individual neuron highly correlated with a specific factual knowledge, capable of influencing the corresponding knowledge expression. QR Cluster represents a coarse grouping of neurons associated with a specific fact. This cluster may include noisy neurons and require further refinement. Common Neuron is consistently activated by a wide range of inputs, representing general knowledge or concepts.
The overall framework is shown in Figure 1. Our framework first resorts to the proxy task of Multi-Choice QA to deal with long-form texts. Starting with a given input, the framework employs Neuron Attribution to derive a QR Cluster. Each neuron within this cluster is assigned an attribution score that indicates its relevance to the query. Next, we apply Inverse Cluster Attribution to attenuate the influence of neurons that frequently occur across multiple clusters. Finally, we identify Common Neurons, as those lacking discriminative power in determining query relevance and representing common knowledge or concepts. Refining the extraction of QR neurons by excluding these common neurons enhances the precision in identifying critical neural correlates.
@@ 4.1 Multi-Choice QA Transformation
Multi-choice QA is widely used in a variety of real-world educational assessments and standardized tests. Meanwhile, many known benchmarks such as MMLU (Hendrycks et al. 2020) and Big-bench (Srivastava et al. 2023) use multi-choice QA to evaluate the breadth and depth of a model’s knowledge. Therefore, we adopt the proxy task of multi-choice QA to study the knowledge associations in LLMs. To deal with free-form answers, we advocate for the transformation of questions and their corresponding answers into a multiple-choice framework, as illustrated in Figure 1. This approach involves the generation of distracted options by randomly sampling answers within the same domain. Following this, the LLM is prompted to produce only the option letter. Subsequently, we investigate the neurons correlated with the input. To mitigate the impact of randomness, we devise multiple prompt templates and systematically shuffle the order of options to prevent the model from learning spurious correlations based on option letters. These prompt templates are detailed in Table A3 in the Supplementary Material in the extended version of this paper [^1] https://arxiv.org/abs/2406.10868 (SM in short in the remainder of this paper).
@@ 4.2 Neuron Attribution
Here, the numerator means that we sum up the scores of different instantiated templates together as the initial attribution score. The denominator $Z$ is the normalization factor obtained by summing the initial attribution scores of all neurons. Since the number of prompts for each query may vary and the initial attribution scores may be scaled differently, we use normalization to make the attribution scores comparable across queries.
@@ 4.4 Common Neurons
We observe that some neurons with a relatively high attribution score are still shared across clusters. Through case studies (as shown in Table 4), we demonstrate that they express commonly used concepts such as option letters (“A” and “B”) or stop words (“and” and “the”). Therefore, we count the frequency of each neuron across clusters. If the frequency is higher than the $u\%$ of total clusters, we assign the given neuron into the common neuron set.
@@ 4.5 Obtaining QR Neurons
Given a query, the final score of a neuron is given by:
@@ 5.1 Experimental Settings
The statistics of our datasets are shown in Table 2 and examples can be found in Table A4 in the SM.
@@
We compare QRNCA to other neuron-level baselines [^2] We do not compare to ROME (Meng et al. 2022a) since it locates layers instead of neurons. Also, we do not compare to task-specific methods.: Random Neuron are randomly selected from FFNs, making sure they have the same number of neurons as QRNCA; Activation selects neurons with high activated values. Kowledge Neuron is adapted from knowledge attribution (Dai et al. 2022) by using the multi-choice QA task; QRNCA wo/ ICA only uses neuron attribution (Eq 5) to obtain relevant neurons, which dose not involve the computation of Inverse Cluster Attribution; QRNCA w/ Common Neuron is a variant without removing common neurons.
@@
Table 3: Comparisons of different knowledge locating methods for Llama-2-7B. The metric here is the Probability Change Ratio (PCR) described in Section 5.1. Details are shown in Table A2 in the SM.
@@
As for the hyper-parameters, the number of estimation steps was set to $m\!=\!16$ and the attribution threshold $t$ to 0.2 times the maximum attribution score. The template number was $|\mathcal{Q}|\!=\!3$, the frequency $u$ for obtaining common neurons was 30%, and the top-$v$ for select coarse neurons was 20. We ran all experiments on three NVIDIA-V100. It took 120 seconds on average to locate neurons for a query with three prompt templates. For each domain and language, the average number of detected QR neurons is between 12 and 17 (see Table A1 in the SM). Hyper-parameters are selected based on a hold-out set of biology queries with 50 samples.
@@ 5.3 QR Neurons Can Impact the Knowledge Expression
Table 3 presents the overall performance of various methods. Our QRNCA method consistently outperforms other baselines, evidenced by its higher PCR. This indicates that our identified QR neurons significantly affect the probability of correct answers while exerting a relatively low impact on unrelated queries. For instance, our method achieves a boosting ratio of 41.2 on the language dataset, the highest among the baselines. Additionally, both our proposed ICA and the removal of common neurons provide further benefits in locating neurons, as evidenced by the worse performance of the two QRNCA variants.
Furthermore, Figure 3 illustrates the percentage change in probability for each domain and language after boosting neuron values. Again, we can clearly observe the effectiveness of our detected QR neurons. Additionally, we performed experiments on Mistral-7B. The results, presented in Figure A4 in the SM, consistently support our conclusions.
@@ 7 Conclusion
In this study, we introduce a novel framework, QRNCA, for identifying neurons in LLMs for long-form answers, extending beyond triplet facts. To validate our approach, we curate two datasets encompassing diverse domains and languages. Our experimental results show that our method outperforms existing baselines in identifying associated neurons. Additionally, this study pioneers the exploration of localized knowledge regions in LLMs and demonstrates Llama contains knowledge-specific regions in the middle layers while language-specific neurons tend to be distributed across different layers. Further, we prototype two potential usages of identified neurons in applications such as knowledge editing and neuron-based prediction. We hope that our findings are beneficial for further research in understanding the knowledge mechanisms underlying LLMs.
Acknowledgments
This research was partially supported by the UKRI INDICATE project (Grant No. EP/Y017749/1), by the ERC under the EU’s Horizon 2020 research and innovation program (grant agreement No. 101020934, ADIX), and by J.P. Morgan and the Royal Academy of Engineering under the Research Chairs and Senior Research Fellowships scheme.